OpenStudy (anonymous):

A truck is rented at $40 per day plus a charge per mile of use. The truck traveled 15 miles in one day and the total charge was $115. Which equation can be used to find the charge x, in dollars, per mile traveled? 40x + 15 = 115 40 + 15x = 115 40x + 15x = 115 40 + 15 = 115x

OpenStudy (anonymous):

If charge per mile is "x", then the total charge per day will be: (Fixed Charge=40)+(Number of miles)x=40+15x=115 which is your second choice I am afraid